# The Hofstadter consecutive-sum sequence omits infinitely many positive integers

### 10 Mar 2026 v1

@@ 4.3. Bloom’s convex difference-set bound and the basic recursion
The key additive-combinatorial input is the following difference-set bound for finite convex sets, due to Bloom [1, Theorem 2].
@@ 5. Concluding remarks
For the classical Hofstadter consecutive-sum sequence (1.1), the present paper provides both qualitative and quantitative progress on Problem 1.1. On the qualitative side, Corollary 1.5 shows that
$a_{n}=n+\omega(1),$
so the sequence omits infinitely many positive integers. On the quantitative side, Theorem 1.6 proves that for every $\varepsilon>0$,
$a_{n}\ll_{\varepsilon}n^{4175/2506+\varepsilon}.$
In particular, the sequence (1.1) has polynomial growth.
There remains, however, a substantial gap between these two conclusions, and the upper bound above is unlikely to be optimal. It is worth noting that, assuming the conjectural near-quadratic lower bound $|A-A|\gg|A|^{2-o(1)}$ for every finite convex set AA—a conjecture often attributed to Erdős and Hegyvári (see [3, p. 2])—the method of Section 4 would yield $a_{n}\leq n^{1+o(1)}$. However, even under this conjectural input, the present iteration argument still does not seem to give the stronger estimate $a_{n}\ll n$.
The available numerical evidence suggests that the true growth is much closer to linear. Writing
@@
it is natural to ask for the true order of magnitude of $b_{n}$. To this end, we computed $b_{n}$ for $1\leq n\leq 30000$ and plotted the normalized quantities $b_{n}/n^{1/k}$ for $k=5,4,3,2$; see Figure 1. The ratio $b_{n}/n^{1/2}$ appears to decrease with $n$, whereas $b_{n}/n^{1/5}$ appears to increase, which is consistent with a power-law heuristic of the form
@@
for some exponent $\alpha$ satisfying $1/5<\alpha<1/2$. Among these plots, the curve $b_{n}/n^{1/3}$ appears comparatively flat on the computed range, suggesting that $\alpha$ might be close to $1/3$, although it is not clear how representative this behavior is asymptotically.
Figure 1. Numerical data for sequence (1.1): the top panel shows $b_{n}=a_{n}-n$ (on a logarithmic $n$-axis), and the remaining panels plot the ratios $b_{n}/n^{1/k}$ for $k=5,4,3,2$ over $1\leq n\leq 30000$.
Acknowledgments. An earlier version of this paper, containing only the result $a_{n}=n+\omega(1)$, appeared as a short note in the comments section of the Erdős Problems website [2]. Matthew Bolan also independently obtained the qualitative result that the classical sequence omits infinitely many positive integers; see again the comments section of the Erdős Problems website [2]. We thank Ingo Althöfer for suggesting the generalization of Conjecture 1.2 to the setting of Definition 1.3. We also thank Boris Alexeev and Wolfram Bernhardt for computing and sharing the numerical data posted in the comments section of the Erdős Problems website [2].
Declaration of AI usage. An AI assistant (ChatGPT, model: GPT-5.2 Thinking) was used only during the brainstorming phase related to Theorem 1.6, and to draft the Python code used to generate Figure 1. Apart from this limited assistance, all mathematical arguments, proofs, and the writing of this paper were carried out by the author.

### 23 Mar 2026 v2

@@ 5.3. Bloom’s convex difference-set bound and the basic recursion
The key additive-combinatorial input is the following difference-set bound for finite convex sets, due to Bloom [3, Theorem 2].
@@ 6. Concluding remarks
6.1. Generalized seeds of size $2$ or larger
Although the body of this paper has focused exclusively on the classical seed $(1,2)$, the same circle of ideas applies much more generally. If one starts from any fixed positive seed
$\mathbf{u}=(u_{1},\dots,u_{s})\qquad(s\geq 2)$
and then follows the same greedy rule, the monotonicity argument of Section 3, together with the eventual-linearity contradiction based on powers of $2$, still yields the analogue of Theorem 1.3. Likewise, the convex-difference-set argument of Section 5 extends, after routine changes in notation, to give a polynomial upper bound of the same shape as in Theorem 1.5. We have chosen not to include the full details here in order to keep the paper focused on Hofstadter’s original sequence (1.1).
6.2. Later upper-bound improvements
After this paper was completed, Sothanaphan pointed out [^1] See the comments section of the Erdős Problems website [5]. that Cushman’s recent bound
$|A-A|\gg_{\varepsilon}|A|^{8/5+1/3440-\varepsilon}$
for finite convex sets can be combined with the argument of Section 5 to yield the improved estimate
$a_{n}\ll_{\varepsilon}n^{688/413+\varepsilon}.$
See [12, 4]. Since this refinement does not alter the underlying method, we have chosen to keep the original exposition.
6.3. Open questions
A substantial gap remains between Theorems 1.4 and 1.5. It seems likely that the true growth of $a_{n}$ is much closer to linear than our present upper bound indicates. Writing
@@ 6.3. Open questions
a natural next step is to determine the true order of magnitude of $b_{n}$. Most notably, one may ask whether $b_{n}$ is sublinear.
Problem 6.1.
Is it true that
$a_{n}=n+o(n)?$
Equivalently, does $b_{n}=o(n)$ hold?
To complement this problem, we computed $b_{n}$ for $1\leq n\leq 30000$ and plotted the normalized quantities $b_{n}/n^{1/k}$ for $k=5,4,3,2$; see Figure 1. The ratio $b_{n}/n^{1/2}$ appears to decrease with $n$, whereas $b_{n}/n^{1/5}$ appears to increase, which is consistent with a power-law heuristic of the form
@@
for some exponent $\alpha$ satisfying $1/5<\alpha<1/2$. Among these plots, the curve $b_{n}/n^{1/3}$ appears comparatively flat on the computed range, suggesting that $\alpha$ might be close to $1/3$, although the available data are far from decisive.
Figure 1. Numerical data for the sequence (1.1): the top panel shows $b_{n}=a_{n}-n$ (on a logarithmic $n$-axis), and the remaining panels plot the ratios $b_{n}/n^{1/k}$ for $k=5,4,3,2$ over $1\leq n\leq 30000$.
Acknowledgments. An earlier version of this paper, containing only the qualitative conclusion $a_{n}=n+\omega(1)$, appeared as a short note in the comments section of the Erdős Problems website [5]. Matthew Bolan also independently obtained the qualitative result that the classical sequence omits infinitely many positive integers; see again the comments section of the Erdős Problems website [5]. We are grateful to Ingo Althöfer for encouraging us to think about extensions to more general initial seeds. We thank Yann Bugeaud for pointing out that the method in our proof of $a_{n}=n+\omega(1)$ can in fact be strengthened to yield the bound of Theorem 1.4. We also thank Norbert Hegyvári for sharing the papers [1, 9, 13] on consecutive sums. Finally, we thank Boris Alexeev and Wolfram Bernhardt for computing and sharing the numerical data posted in the comments section of the Erdős Problems website [5].
Declaration of AI usage. An AI assistant was used only during the brainstorming phase related to Theorem 1.5, and to draft the Python code used to generate Figure 1. Apart from this limited assistance, all mathematical arguments, proofs, and the writing of this paper were carried out by the author.